# A General Framework for Joint Multi-State Models

### v2

@@ Abstract
Classical joint modeling approaches often rely on competing risks or recurrent event formulations to describe complex processes involving evolving longitudinal biomarkers and discrete event occurrences, but these frameworks typically capture only limited aspects of the underlying event dynamics.
We propose a general multi-state joint modeling framework that unifies longitudinal biomarker dynamics with multi-state time-to-event processes defined on arbitrary directed graphs. The proposed framework accommodates arbitrary directed transition graphs, nonlinear longitudinal submodels, and scalable inference via stochastic gradient descent. This formulation encompasses both Markovian and semi-Markovian transition structures, allowing recurrent cycles and terminal absorptions to be naturally represented. The longitudinal and event processes are linked through shared latent structures within nonlinear mixed-effects models, extending classical joint modeling formulations.
We derive the complete likelihood, establish conditions for identifiability, and develop scalable inference procedures based on stochastic gradient descent to enable high-dimensional and large-scale applications. In addition, we formulate a dynamic prediction framework that provides individualized state-transition probabilities and personalized risk assessments along complex event trajectories.
@@ 1 Introduction
Joint modeling of longitudinal and time-to-event data has become an essential tool of modern biostatistics [papageorgiou2019overview], particularly for dynamic prediction in clinical applications. Classical joint models typically couple a longitudinal biomarker process with a single time-to-event outcome [rizopoulos2012joint], allowing for the integration of biological knowledge and individual heterogeneity via shared latent structures. However, many real-world processes involve multiple possible outcomes, intermediate stages, or recurrent events, which cannot be fully captured by a single-event framework [KrolMauguen2017]. In such settings, multi-state models provide a more flexible approach [ferrer2016joint].
Multi-state models represent an overall joint process of clinical course as a series of discrete stages or health states that occur sequentially [lovblom2024modeling]. In biostatistics, these models are widely used for survival and reliability analysis, allowing for a richer and more accurate representation by capturing alternative paths to an event of interest, intermediate events, and progressive disease. Key components of multi-state models include transition intensity functions, which denote the instantaneous risk of moving from one state to another, and transition probability functions, which describe the probability of transition over longer intervals. Often, these models assume the Markov property, where future transitions depend only on the current state, simplifying the transition intensity functions [asanjarani2022estimation].
The Markov assumption can be relaxed by adopting a semi-Markov formulation, in which the transition probabilities depend not only on the current state but also on the sojourn time, effectively resetting the time scale after each transition.
The link between multi-state models and joint models arises when a multi-state process is integrated as a component within a broader joint modeling framework . While multi-state models, such as those described in the “sequential state framework” primarily focus on movements between discrete states, joint models operate under a “parallel trajectory framework” that combines a longitudinal process with a time-to-event process.
@@
This paper is organized as follows. Section 2 reviews the background on multi-state and joint modeling frameworks. Section 3.2 introduces the proposed general multi-state joint modeling framework, detailing the likelihood formulation and identifiability conditions. Section 4 describes the stochastic-gradient–based inference algorithm, while Section 5 presents the dynamic prediction framework. Section 6 reports simulation experiments that assess parameter recovery, identifiability, and convergence properties of the proposed model, and Section 7 applies the framework to data from the PAQUID cohort. Finally, Section 8 concludes with perspectives and future research directions.
@@ 2 Background
Joint modeling of longitudinal and time-to-event data provides a unified statistical framework to analyze the interplay between continuous biomarker dynamics and event occurrence processes. In its classical form, this framework couples a mixed-effects model describing individual biomarker trajectories with a survival model for event times [rizopoulos2012joint]. Such models have become central in biomedical research, particularly for dynamic prediction and personalized risk assessment [papageorgiou2019overview]. However, these traditional joint models generally focus on a single terminal event, limiting their ability to represent more complex event histories that include intermediate, recurrent, or competing events. To address these limitations, multi-state extensions have been developed, providing a natural framework for describing transitions between multiple clinical or biological states over time [ferrer2016joint, KrolMauguen2017].
@@ 3 A General Multi-State Joint Modeling Framework
Classical joint models typically link a longitudinal biomarker process to a single time-to-event outcome. While effective for terminal events, this design cannot represent more complex disease trajectories involving intermediate or recurrent events. To address these limitations, we extend the nonlinear joint modeling framework to incorporate a multi-state process, thereby enabling a unified analysis of longitudinal biomarkers and discrete transitions between multiple health states.
@@ 3.1 Notation Conventions
The latent random effects $b_{i}\sim\mathcal{N}(0,Q)$ capture individual heterogeneity, while $\psi_{i}=f(\gamma,X_{i},b_{i})$ denotes the subject-specific parameters in the longitudinal submodel. The filtration $\mathcal{H}_{i}(t)$ represents the observed history of the longitudinal process up to time $t$.
Transition times and states are denoted by $(T_{i\ell},S_{i\ell})$, with right-censoring time $C_{i}$. We also define the index of the last observed transition before time $t$ denoted by $m_{i}(t)\coloneq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$.
@@ 3.2 Extending Multi-State Joint Models
Joint models traditionally focus on a single time-to-event outcome, which limits their ability to represent complex disease trajectories involving multiple intermediate and recurrent events. To address this limitation, we extend the classical nonlinear joint modeling framework to incorporate multi-state processes, enabling the joint analysis of longitudinal biomarkers and transitions between multiple health states. This is achieved by introducing a directed graph structure encoding all permissible transitions and supports both Markov and semi-Markov assumptions.
@@ 3.2.3 Transition Intensities and Survival Functions
$\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})=\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i,\ell+1}\leq t+\delta,S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1}>t,T_{i\ell}=t_{0},S_{i\ell}=k)}{\delta}.$
The cumulative intensity and the corresponding survival function are
@@ 3.3 Multi-State Joint Model Specification
Here, $h$ and $f$ define the nonlinear mixed-effects submodel, $g^{k^{\prime}\mid k}$ represents the link between biomarker dynamics and the transition intensity, and $\lambda_{0}^{k^{\prime}\mid k}$ denotes the baseline hazard associated with transition $(k\to k^{\prime})$. This structure extends the classical joint model to arbitrary directed graphs and to both Markovian and semi-Markov specifications. The coefficient $\alpha^{k^{\prime}\mid k}$ quantifies the effect of the longitudinal biomarker dynamics on the instantaneous risk of transitioning from state $k$ to state $k^{\prime}$. More concretely, for a one-unit increase in $g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})$ (holding other covariates constant), the hazard ratio for transition $(k\to k^{\prime})$ is $\exp(\alpha^{k^{\prime}\mid k})$. Likewise, $\beta^{k^{\prime}\mid k}$ represents the effect of baseline (or time-varying) covariates $X_{i}$ on that same transition: a one-unit increase in a covariate $X_{ij}$ multiplies the hazard by $\exp(\beta^{k^{\prime}\mid k}_{j})$. Thus, $\alpha^{k^{\prime}\mid k}$ captures the degree to which the biomarker trajectory (via its latent parameter $\psi_{i}$ or function $g^{k^{\prime}\mid k}$) influences the transition risk, while $\beta^{k^{\prime}\mid k}$ captures the direct effect of covariates on that transition’s risk, beyond the biomarker pathway.
@@ 3.4 Model Selection and Information Criteria
$\mathrm{AIC}=-2\log\mathcal{L}_{\text{marginal}}(\hat{\theta};X,Y,\mathcal{T},C)+2k,$
@@ 4 Statistical Inference
Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [kuhn2004coupling], Laplace approximation [wolfinger1993laplace], Gauss-Hermite quadrature, stochastic gradient ascent with Robbins–Monro updates [robbins1951stochastic], and MCMC-based approximations such as Metropolis-Hastings and Hamiltonian Monte Carlo. These strategies are implemented in software such as JMBayes [jmbayes22024, rizopoulos2020package].
Another approach requiring mild regularity assumptions on the $\log$ marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [caillebotte2025estimation, baey2023efficient] using Fisher’s identity and following the Robbins-Monro procedure [robbins1951stochastic].
Under interchangeability of integral and differentiation, setting $x=(X,Y,\mathcal{T},C)$ for convenience, the Fisher identity writes
@@ 5 Dynamic prediction
$\hat{\chi}_{i}=\mathbb{E}_{\mathcal{T}^{\prime}_{i}\sim\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t)}\left(\chi^{*}(\mathcal{T}^{\prime}_{i})\right)\approx\frac{1}{B}\sum_{k=1}^{B}\chi^{*}(\mathcal{T}_{i}^{(k)}),\,\mathcal{T}_{i}^{(k)}\overset{\text{i.i.d.}}{\sim}\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t).$
@@
Assume there exists $\chi:\bigcup_{n\geq 1}(\mathbb{R}\times\mathbb{S})^{n}\to\mathbb{Y}$ and $\tau_{i}$ a stopping time for the filtration $\mathcal{F}_{in}=\sigma\left((T_{il},S_{il})_{0\leq l\leq n}\right)$ such that:
@@
$\operatorname{depth}(G)=\max_{(v_{0},v_{1},\dots,v_{k})}k,$
@@ 6 Simulation Study
The purpose of this simulation study is to evaluate the finite-sample performance, identifiability, and convergence properties of the proposed inference algorithm under controlled conditions where the true parameters are known. We simulate data from a three-state semi-Markov process coupled with a nonlinear longitudinal biomarker trajectory, representing a simplified disease progression model. This setup allows us to assess both statistical accuracy and computational scalability of the stochastic-gradient estimation procedure.
@@ 7.2.1 Inference
Figure 8: Inferred parameters and their estimated standard deviations. Bold values mean $p<0.05$.

### v3

@@ Abstract
Conventional joint modeling approaches generally characterize the relationship between longitudinal biomarkers and discrete event occurrences within terminal, recurring or competing risk settings, thereby offering a limited representation of complex, multi-state trajectories.
We propose a general multi-state joint modeling framework that unifies longitudinal biomarker dynamics with multi-state time-to-event processes defined on arbitrary directed graphs. The proposed framework also accomodates nonlinear longitudinal submodels and scalable inference via stochastic gradient descent. This formulation encompasses both Markovian and semi-Markovian transition structures, allowing recurrent cycles and terminal absorptions to be naturally represented. The longitudinal and event processes are linked through shared latent structures within nonlinear mixed-effects models, extending classical joint modeling formulations.
We derive the complete likelihood, model selection criteria, and develop scalable inference procedures based on stochastic gradient descent to enable high-dimensional and large-scale applications. In addition, we formulate a dynamic prediction framework that provides individualized state-transition probabilities and personalized risk assessments along complex event trajectories.
@@ 1 Introduction
Joint modeling of longitudinal and time-to-event data has become an essential tool of modern biostatistics [papageorgiou2019overview], particularly for dynamic prediction in clinical applications. Classical joint models typically couple a longitudinal biomarker process with a single time-to-event outcome [rizopoulos2012joint], allowing for the integration of biological knowledge and individual heterogeneity via shared latent structures. However, many real-world processes involve multiple possible outcomes, intermediate stages, or recurrent events, which cannot be fully captured by a single-event framework [KrolMauguen2017]. To address these limitations, specialized frameworks have been developed to handle competing risks and recurrent events. However, these remain constrained to specific event types and fail to accommodate more general multi-state transitions. In such settings, multi-state models provide a more general and flexible approach [ferrer2016joint].
Multi-state models represent an overall joint process of clinical course as a series of discrete stages or health states that occur sequentially [lovblom2024modeling]. In biostatistics, these models are widely used for survival and reliability analysis, allowing for a richer and more accurate representation by capturing alternative paths to an event of interest, intermediate events, and progressive disease. Key components of multi-state models include transition intensity functions, which denote the instantaneous risk of moving from one state to another, and transition probability functions, which describe the probability of transition over longer intervals. Often, these models assume the Markov property, where future transitions depend only on the current state, simplifying the transition intensity functions [asanjarani2022estimation]. This assumption can be relaxed by adopting a semi-Markov formulation, in which the transition probabilities depend not only on the current state but also on the sojourn time, effectively resetting the time scale after each transition.
The link between multi-state models and joint models arises when a multi-state process is integrated as a component within a broader joint modeling framework . While multi-state models, such as those described in the “sequential state framework” primarily focus on movements between discrete states, joint models operate under a “parallel trajectory framework” that combines a longitudinal process with an acyclic time-to-event process.
@@
This paper is organized as follows. Section 2 reviews the background on multi-state and joint modeling frameworks. Section 3.2 introduces the proposed general multi-state joint modeling framework, detailing the likelihood formulation and proposed model selection creteria. Section 4 describes the stochastic-gradient–based inference algorithm, while Section 5 presents the dynamic prediction framework. Section 6 reports simulation experiments that assess parameter recovery and convergence properties of the proposed model, whereas Section 7 applies the framework to data from the PAQUID cohort. Finally, Section 8 concludes with perspectives and future research directions.
@@ 2 Background
Joint modeling of longitudinal and time-to-event data provides a unified statistical framework to analyze the interplay between continuous biomarker dynamics and event occurrence processes. In its classical form, this framework couples a mixed-effects model describing individual biomarker trajectories with a survival model for event times [rizopoulos2012joint]. Such models have become central in biomedical research, particularly for dynamic prediction and personalized risk assessment [papageorgiou2019overview]. However, these traditional joint models generally focus either on a single terminal event, competing, or recurrent events, limiting their ability to represent more complex event histories that include intermediate, recurrent, or competing events. To address these limitations, multi-state extensions have been developed, providing a natural framework for describing transitions between multiple clinical or biological states over time [ferrer2016joint, KrolMauguen2017].
@@ 3 A General Multi-State Joint Modeling Framework
To address the limitations imposed by traditional joint modeling frameworks, we extend them to incorporate a multi-state process, thereby enabling a unified analysis of longitudinal biomarkers and discrete transitions between multiple health states.
@@ 3.1 Notation Conventions
The latent random effects $b_{i}\sim\mathcal{N}(0,Q)$ capture individual heterogeneity, while $\psi_{i}\coloneqq f(\gamma,X_{i},b_{i})$ denotes the subject-specific parameters in the longitudinal submodel. The filtration $\mathcal{H}_{i}(t)$ represents the observed history of the longitudinal process up to time $t$.
Transition times and states are denoted by $(T_{i\ell},S_{i\ell})$, with right-censoring time $C_{i}$. We also define the index of the last observed transition before time $t$ denoted by $m_{i}(t)\coloneqq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$.
@@ 3.2 Extending Multi-State Joint Models
The joint model proposed by Ferrer et al. [ferrer2016joint] established a foundational link between longitudinal biomarkers and multi-state processes through shared random effects. However, their formalism relies on binary censoring indicators to characterize transitions, thereby inherently restricting the modeling framework to Directed Acyclic Graphs (DAGs) and excluding recurrent events. To overcome this limitation, we present a new method based on arbitrary graph transitions. By shifting away from binary indicators toward a more flexible graph-theoretic specification, our approach accommodates complex disease trajectories, including recurrent events and cyclic transitions that cannot be captured within a DAG-based structure.
@@ 3.2.3 Transition Intensities and Survival Functions
$\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})\coloneqq\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i,\ell+1}\leq t+\delta,S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1}>t,T_{i\ell}=t_{0},S_{i\ell}=k)}{\delta}.$
The cumulative risk and the corresponding survival function are
@@ 3.3 Multi-State Joint Model Specification
Here, $h$ and $f$ define the nonlinear mixed-effects submodel, $g^{k^{\prime}\mid k}$ represents the link between biomarker dynamics and the transition intensity, and $\lambda_{0}^{k^{\prime}\mid k}$ denotes the baseline hazard associated with transition $(k\to k^{\prime})$. This structure extends the classical joint model to arbitrary directed graphs and to both Markovian and semi-Markov specifications.
The coefficient $\alpha^{k^{\prime}\mid k}$ quantifies the effect of the longitudinal biomarker dynamics on the instantaneous risk of transitioning from state $k$ to state $k^{\prime}$. More concretely, for a one-unit increase in $g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})$ (holding other covariates constant), the hazard ratio for transition $(k\to k^{\prime})$ is $\exp(\alpha^{k^{\prime}\mid k})$. Likewise, $\beta^{k^{\prime}\mid k}$ represents the effect of baseline (or time-varying) covariates $X_{i}$ on that same transition: a one-unit increase in a covariate $X_{ij}$ multiplies the hazard by $\exp(\beta^{k^{\prime}\mid k}_{j})$. Thus, $\alpha^{k^{\prime}\mid k}$ captures the degree to which the biomarker trajectory (via its latent parameter $\psi_{i}$ or function $g^{k^{\prime}\mid k}$) influences the transition risk, while $\beta^{k^{\prime}\mid k}$ captures the direct effect of covariates on that transition’s risk, beyond the biomarker pathway.
@@ 3.4 Model Selection and Information Criteria
$\mathrm{AIC}\coloneqq-2\log\mathcal{L}_{\text{marginal}}(\hat{\theta};X,Y,\mathcal{T},C)+2k,$
@@ 4 Statistical Inference
Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [kuhn2004coupling], Laplace approximation [wolfinger1993laplace] and Gauss–Hermite quadrature. These strategies are implemented in software such as JMBayes [jmbayes22024, rizopoulos2020package].
Another approach requiring mild regularity assumptions on the $\log$ marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [caillebotte2025estimation, baey2023efficient] using Fisher’s identity and following the Robbins-Monro procedure [robbins1951stochastic], alongside with MCMC-based approximations such as Metropolis-Hastings and Hamiltonian Monte Carlo.
Under interchangeability of integral and differentiation, setting $x\coloneqq(X,Y,\mathcal{T},C)$ for convenience, the Fisher identity writes
@@ 5 Dynamic prediction
$\hat{\chi}_{i}=\coloneqq\mathbb{E}_{\mathcal{T}^{\prime}_{i}\sim\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t)}\left(\chi^{*}(\mathcal{T}^{\prime}_{i})\right)\approx\frac{1}{B}\sum_{k=1}^{B}\chi^{*}(\mathcal{T}_{i}^{(k)}),\,\mathcal{T}_{i}^{(k)}\overset{\text{i.i.d.}}{\sim}\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t).$
@@
Assume there exists $\chi:\bigcup_{n\geq 1}(\mathbb{R}\times\mathbb{S})^{n}\to\mathbb{Y}$ and $\tau_{i}$ a stopping time for the filtration $\mathcal{F}_{in}\coloneqq\sigma\left((T_{il},S_{il})_{0\leq l\leq n}\right)$ such that:
@@
$\operatorname{depth}(G)\coloneqq\max_{(v_{0},v_{1},\dots,v_{k})}k,$
@@ 6 Simulation Study
The purpose of this simulation study is to evaluate the finite-sample performance and convergence properties of the proposed inference algorithm under controlled conditions where the true parameters are known. We simulate data from a three-state semi-Markov process coupled with a nonlinear longitudinal biomarker trajectory, representing a simplified disease progression model. This setup allows us to assess both statistical accuracy and computational scalability of the stochastic-gradient estimation procedure.
@@ 7.2.1 Inference
Figure 8: Inferred parameters and their estimated standard deviations. Bold values mean $p<0.05$ (variance parameters excluded).